OpenStudy (anonymous):

X plus 4 all over 3 = 2 ?

hartnn (hartnn):

\[\frac{x+4}{3}=2\] start by multiplying 3 on both sides. what u get?

OpenStudy (anonymous):

multiple 3(4) and 3(2) ?

hartnn (hartnn):

no,like this: \[3\frac{x+4}{3}=3*2\] so in left side only x+4 remains ok?

OpenStudy (anonymous):

so would x=2?

hartnn (hartnn):

yup,correct :)